# TREX: Automating LLM Fine-tuning via Agent-Driven Tree-based Exploration

### 15 Apr 2026 v1

@@ 5.1 Experimental Setup
Tasks. We evaluate the proposed TREX framework across 10 tasks from FT-Bench. Each task is fed to the Researcher via a structured template, with the input comprising the task description, evaluation protocol and entry, raw training data (optional, available only for specific tasks), and experimental configurations (see Figure 1 for an example). To ensure time and computational efficiency, we uniformly employ the lightweight model Qwen3-1.7B as the base model for autonomous fine-tuning and cap the maximum number of training samples at 50,000 per experiment.
System Settings. For the Researcher agent, we verify its capability of conducting autonomous, multi-round experimental design across varying backend LLMs, including an open-source model (Qwen3-Next-80B-Thinking) and a proprietary model (Gemini 3 Pro). For the Executor agent, we adopt Claude 4.5 Sonnet, following OpenHands’ recommendations, to maximize the robustness of experiment execution.
@@
Where the $\mathcal{E}_{T}$ is the task-specific evaluation function, while $M_{Base}$ and $M_{FT}$ represent the models before and after autonomous fine-tuning, respectively. $M_{Ref}$ is a superior reference model, and the performance gap between it and the base model provides a baseline for normalizing and quantifying the relative gains achieved. We use Qwen3-235B-2507 as $M_{Ref}$ in experiments.
@@ 5.2 Main results
Figure 3: Comparative performance trajectories of iterative experimental refinement using distinct LLM engines for the task TOMG-Bench and OpenFinData. The left panel illustrates the score progression with Gemini3-Pro-Thinking as the Researcher, while the right panel depicts the corresponding progression with Qwen3-Next-80B-Thinking.
@@ 5.3 Ablation Studies
Figure 4: Performance of TREX across search strategies on oMeBench and GTA tasks. The dashed line represents the score trajectory, while the solid line with $\uparrow$ represents the growth curve of the score trajectory.
@@
Figure 5: Performance of TREX with versus without AIDP tools on oMeBench and GTA tasks. The dashed line represents the score trajectory, while the solid line with $\uparrow$ represents the growth curve of the score trajectory.
@@
Figure 6: Performance of TREX with accessible versus inaccessible bad cases on oMeBench and GTA tasks. The dashed line represents the score trajectory, while the solid line with $\uparrow$ represents the growth curve of the score trajectory.
@@ A.2 AIDP Package
A.3 Experiment Plan Examples

### 22 Apr 2026 v2

@@ 5.1 Experimental Setup
Tasks. We evaluate the proposed TREX framework across 10 tasks from FT-Bench. Each task is fed to the Researcher via a structured template, with the input comprising the task description, evaluation protocol and entry, raw training data (optional, available only for specific tasks), and experimental configurations (see Figure 1 for an example). To ensure time and computational efficiency, we uniformly employ the lightweight model Qwen3-1.7B [qwen3technicalreport] as the base model for autonomous fine-tuning and cap the maximum number of training samples at 50,000 per experiment.
System Settings. For the Researcher agent, we verify its capability of conducting autonomous, multi-round experimental design across varying backend LLMs, including an open-source model (Qwen3-Next-80B-Thinking [qwen3technicalreport]) and a proprietary model (Gemini 3 Pro [googledeepmind2025gemini3pro]). For the Executor agent, we adopt Claude 4.5 Sonnet, following OpenHands’ recommendations, to maximize the robustness of experiment execution.
@@
Where the $\mathcal{E}_{T}$ is the task-specific evaluation function, while $M_{Base}$ and $M_{FT}$ represent the models before and after autonomous fine-tuning, respectively. $M_{Ref}$ is a superior reference model, and the performance gap between it and the base model provides a baseline for normalizing and quantifying the relative gains achieved. We use Qwen3-235B-2507 [qwen3technicalreport] as $M_{Ref}$ in experiments.
@@ 5.2 Main results
In Appendix A.3, we provide the trajectories of performance scores across all experiments throughout the iterative process.
@@
Figure 3: Comparative performance trajectories of iterative experimental refinement using distinct LLM engines for the task TOMG-Bench and OpenFinData. The left panel illustrates the score progression with Gemini 3 Pro as the Researcher, while the right panel depicts the corresponding progression with Qwen3-Next-80B-Thinking.
@@ 5.3 Ablation Studies
Figure 4: Performance of TREX across search strategies on oMeBench and GTA tasks. The dashed line represents the score trajectory, while the solid line with $\uparrow$ represents the frontier of the score trajectory.
@@
Figure 5: Performance of TREX with versus without AIDP tools on oMeBench and GTA tasks. The dashed line represents the score trajectory, while the solid line with $\uparrow$ represents the frontier of the score trajectory.
@@
Figure 6: Performance of TREX with accessible versus inaccessible bad cases on oMeBench and GTA tasks. The dashed line represents the score trajectory, while the solid line with $\uparrow$ represents the frontier of the score trajectory.
@@ A.2 AIDP Package
A.3 Additional Experiment Results
Figure 7 and Figure 8 illustrate the performance trajectories of TREX across all tasks in FT-Bench.
Figure 7: Comparative performance trajectories of iterative experimental refinement using distinct LLM engines for tasks oMeBench, ACI-Bench, GTA, and HOC. The left panel illustrates the score progression with Gemini 3 pro as the Researcher, while the right panel depicts the corresponding progression with Qwen3-next-80b-thinking.
Figure 8: Comparative performance trajectories of iterative experimental refinement using distinct LLM engines for the task LawBench and EconlogicQA, SST-2 and CS-Bench. The left panel illustrates the score progression with Gemini 3 pro as the Researcher, while the right panel depicts the corresponding progression with Qwen3-next-80b-thinking.
A.4 Experiment Plan Examples